# Nonlinear phase synchronization and the role of spacing in shell models

### v1

@@ IV.1 Shell spacing dependence
Helical shell models provide a rich framework, in which a set of interacting triads are connected in a way that mirrors the actual interactions among two sequences of helical modes in three dimensional Navier-Stokes turbulence. Here, we consider the phase coherence of these two sequences, as a function of the inter-shell ratio $g$, in order to investigate differences in syncronizability and to compare with the single chain GOY model.
Due to their more complicated structure, helical shell model are have been less explored in literature. Here, we restrict our analysis to shell models constructed from a single class of interactions, as in Refs. [59, 60], focusing first on those classes that can sustain a forward energy cascade for local interactions, namely SM1, SM2 and SM3. We note that:
  • SM1 behaves exactly as two uncoupled GOY models. As a result, its synchronization properties are equivalent to those of a single chain as has been characterized in detail in the first part of this work, showing a the transition from the standard intermittency ($g=2$) to strong intermittency as the inter-shell spacing $g$ is decreased.
@@
  • In stark contrast to all the examples that we have seen earlier, SM3 presents a level of intermittency comparable to SM1 for $g=2$ but the intermittency corrections decrease drastically in the small $g$ limit, and the model exhibits monofractal scaling [54].
Those behaviors are presented for clarity in Fig 12.
Figure 12: Plot of the Scaling exponent $\xi_{p}$of the $p-$th order structure function as a function of $p$, for different Helical shell models corresponding to the classes SM1, SM2, and SM3 (indicated by different colors). The different markers styles represent the two inter-shell spacings $g=2$ and $g=1.2$.
The time behavior of the SM3 model for the two representative cases with inter-shell ratio $g=1.2$ and $g=2$ is presented in Fig. 13 showing the evolution of the compensated energy spectrum.
In intermittent case ($g=2$), we observe bursty fluctuations around the global scaling $k_{n}^{-2/3}$, less defined but similar to those seen for the GOY model evolution. Although no clear signature of instantons is to be seen here. It is worth noting that, instantons have been observed in helical shell models when initialized from a zero-background rather than in conditions of a steady state energy cascade. In particular for SM3 class, structures that are less steep than the Kolmogorov scaling, are reported in Ref. [61].
In contrast, in the non-intermittent case $(g=1.2)$, while some local coherent events, that are carrying energy to smaller scales are still present, they do not seem to propagate across many shells as in the intermittent case possibly because the compensated spectra look more noisy at smaller scales. This suggests that triadic phases lose their ability to self-organize in a way that accommodate strong flux events involving a large portion of the inertial range.
@@
Figure 13: Temporal evolution of compensated energy spectrum, $log(k_{n}^{2/3}E_{n}$) of different shells (index $n$ in the $y$-axis) shown for the two values of spacing $g=2$ and $g=1.2$ considered (top to bottom). The plots illustrate the reduction of coherent structures in energy cascade as $g$ decreases.
@@
Figure 14: Normalized PDFs of the global Kuramoto parameter $R$, and the associated average phase $\Phi$ are shown in the top left and right panels respectively, in two representative cases $g=1.2$ and $g=2$. The bottom left and right panels show respectively the PDFs of the local Kuramoro parameter $R_{n}$ and the associated average phase $\Phi_{n}$ for different shell wave-numbers as indicated in the colorbar. .
@@
The phase coherence in the two regimes is shown in Fig. 14, where normalized PDFs of the global Kuramoto order parameter $R$, and the associated average phase $\Phi$ are presented, for the two inter shell spacing $g=1.2$ and $g=2$. A clear difference in the level of synchronization is observed, as the systems moves from intermittent to non intermittent behavior, which is manifested by a marked reduction in triadic phase coherence. The monofractal case exhibits a narrow distribution at low $R$ values, in contrast to the broader distribution of $g=2$ case, indicating how incoherent events dominate the system dynamics. At the same time, the PDF of the average triadic phase $\Phi$ appears broader as the coherence is lost, even though a clustering around $\pi/2$ can still be observed, sustaining a forward energy flux, on average.
@@
In the bottom panels, we present the PDFs of the parameters that characterize local coherence, namely $R_{n}$ and $\Phi_{n}$, for the shells in the inertial range. The local level of phase coherence appear to be universal, i.e. independent of the scale within the inertial range. Moreover, the $g=2$ case displays a distribution of local coherence significantly different from that of the GOY model (presented in Fig. 7), also due to the differences in the definition of the local Kuramoto parameters in the two cases, even though both models show comparable levels of intermittency.
@@
Notably, local coherence persist despite the loss of global synchronization as we can note comparing the PDFs of $R_{n}$ in the two cases $g=2$ and $g=1.2$. When the inter-shell spacing is decreased the position of the maxima shifts slightly toward lower coherence levels, and only highly local coherent events ($R_{n}\simeq 0,8$) are significantly less probable. In contrast, the statistic of the global coherence $\Phi$ show a marked difference with a predominance of incoherent events for $g=1.2$. This statistical features are consistent with the observed temporal dynamics, characterized not by structures that travels over many shells but by events that involve neighboring scales.
@@
Looking at these results, concerning forward cascades, we conclude that the Kuramoto parameter defined using the triadic phases is a useful order parameter for studying phase coherence and intermittency in a network of triad interactions. However, it is important to stress that it is not a unique indicator of intermittency: for instance for the same inter shell ratio $g=2$ the PDF of the Kuramoto parameter differs between GOY model and the helical model SM3, even though they present very similar levels of intermittency. This suggest that the Kuramoto parameter is also sensible to the topological structure of the network of interacting triads. Nonetheless once the topology is fixed it proves to be a highly sensible proxy of the modification of anomalous scaling behavior, both when increasing and decreasing the level of intermittency.
@@ V Inverse cascades
Despite the success of shell models in modeling forward cascades, attempts to extend them, in order to reproduce the inverse energy cascade observed in 2D turbulence - simply by changing the dimensionality of the conserved invariants - have generally failed, mainly because of the tendency of the inverse cascade in these models to be overwhelmed by statistical shell-equipartition solutions (see, e.g., Refs. [62, 39]). To overcome this issue, different generalization has been proposed, such as shell model on hierarchical trees [63, 64], or models based on spiral chains written for energy [24] or a model only for amplitudes, or one based on a stochastic cascade mechanism [65]. These results suggest that the self-consistently evolving phase equation plays a crucial role in suppressing the inverse cascade.
In this work, we will focus on the two shell model that, to our knowledge, are able to present a genuine inverse energy cascade: the elongated helical shell model proposed by De Pietro, Biferale, and Mailybaev [60] and the model studied by Tom and Ray [49] (see also Ref. [45] as a shell model capable of driving an inverse cascade for resonating waves in the context of wave turbulence). Our goal here is to analyze the statistical properties of triadic phases and their implication in the inverse cascade solutions produced by these models as a complementary example to the forward cascade case.
Since inverse cascades are typically non-intermittent and lack of coherent bursts, we do not show a detailed investigation of the synchronizability of those shell models. They are all boringly non-intermittent in that respect. However it is interesting to show the forms of the PDFs of the triadic phases together with the fluxes and the resulting spectra, in order to demonstrate how the triadic phases organize statistically, to drive an inverse cascade even in the presence of strong phase randomness. Hopefully, understanding how triadic phase organization works in such reduced representations as shell models, will provide the key to explaining how the cascade mechanism is sustained in turbulence in general.
@@ V.0.1 Inverse energy cascade in an elongated helical shell model
A particular case of Helical interactions, represented by triads of the type (,+,+)(-,+,+)( - , + , + ) or (+,,)(+,-,-)( + , - , - ), i.e. the class SM2, is known to drive inverse cascade for sufficiently elongated triads. The critical level of elongation at which the direction of energy transfer changes as first predicted by Waleffe [58], occurring at t $p/k=0.278$ (assuming $p<k<q$) , was numerically confirmed with an elongated helical shell model [60], which can apparently, also be justified using arguments based on the conservation of pseudo-invariants [66].
In this section we consider this model, representing a unique example of an inverse cascade driven by non local interactions as our to test bed for studying the phase coherence of elongated triads, which drive the inverse cascade process. We recall that for non local interactions the relevant triadic phases centered in the $n$-th shell can be defined as $u_{n-l}^{s_{1}}u_{n}^{s_{0}}u_{n+m}^{s_{2}}=\chi_{n}e^{i\varphi_{n}}$ and $u_{n-l}^{-s_{1}}u_{n}^{-s_{0}}u_{n+m}^{-s_{2}}=\bar{\chi}_{n}e^{i\bar{\varphi}_{n}}$.
Figure 15: Plot of the energy spectrum as a function of the wave number for a helical shell model driven by small-scale forcing (forcing shells indicated in red). The inset plot on the bottom left shows the spectral flux as a function of scale, corresponding to an inverse energy flux in the inertial range. The inset plot on the top right shows the normalized PDFs of triadic phases $\varphi_{n}$, for the shells in the inertial range. The shell number is represented by different colors labeled by the continuous colorbar.
In Figure (15) we present the numerical results obtained for SM2 (elongated) shell model, i.e. defined by $s_{1}/s_{0}=-1$ and $s_{2}/s_{0}=+1$, with an inter-shell ratio $g=\frac{1+\sqrt{5}}{2}$ and sufficiently elongated triads , specifically $l=3$ and $m=1$, with a small scale forcing. This choice of parameters produces an inverse cascade, as we can see from the spectral energy flux and a clear inertial range, with an energy spectrum consistent with Kolmogorov. The normalized probability distributions of triadic phases, presented in the inset, show a higher probability of events with triadic phases equal to $-\pi/2$, which corresponds to the choice that maximizes an inverse energy flux. The distribution appears to have a roughly sinusoidal shape and seems to be universal, in the sense that it does not seem to show any scaling with shell number $n$.
Comparing these PDFs, with the forward cascade case, the probability of such coherent events is notably less pronounced. This weaker phase coherence is even more pronounced at a global level as shown in Figure (2), where the Kuramoto order parameter reveals a relatively incoherent cascade dynamics, accompanied by a clear flattening of the average phase $\Phi$ distribution.
@@ V.0.2 Inverse energy cascades and equipartition solutions in a local shell model
For the Sabra shell model [17], a detailed phase diagram has been proposed in Ref. [67], and a further investigation of the equipartition and the inverse cascade solution has been carried out by Tom and Ray [49]. Specifically, they considered standard inter-shell ratio $g=2$, by setting $a=1$ and imposing energy conservation through the condition $c=-1-b$. This causes the second positive defined invariant (in the range $-2<b<-1$) to have the form $H=\sum_{n}k_{n}^{\alpha}|u_{n}|^{2}\;$ , with a scaling exponent $\alpha=log_{g}\left(\left|\frac{1}{-1-b}\right|\right)$ Notably, the choice of enstrophy conservation, i.e. $\alpha=2$, corresponds to $b=-5/4$ but as we already know this choice fails to drive an inverse cascade. However, they showed that this shell model sustains an inverse energy cascade for values of $b$ in the range $-2<b<b_{c}=-1.63$, with a spectral slope that approach the characteristic scaling of 2D turbulence in the limit $b\to-2$. We have replicated the numerical results presented in [49], so that we can investigate the role of phase coherence in driving a local inverse cascade.
@@
As shown in Figure (16), the model, when forced at small scales with $b=-1.8$, exhibits a clear inertial range associated with an inverse energy cascade, corresponding to a constant, on average, negative energy flux. The probability distribution of triadic phases, displayed in the top right inset, shows how the triadic phases for different shells organize around the value $-\pi/2,$ which maximizes inverse energy transfer. The distribution follows a sinusoidal shape, i.e odd with respect to $\varphi_{n}\to-\varphi_{n}$, with non universal scale-dependent behavior, indicating that triadic phase organization is weaker, (as marked by noisier and broader curves) going to larger scales (slower time scales). In contrast (not shown), when the model exhibits energy equipartition (i.e. for $b_{c}<b<-1$) triadic phases appear noisy and randomly distributed.
To quantify synchronization at a global level, providing an overall view of phase coherence between forward and inverse cascades, we report in Figure (2), the Kuramoto order parameter and the associated average triadic phase $\Phi$. Notably, we observe that inverse cascades are generally driven by much weaker phase organization. Among the two inverse cascade cases considered, the local inverse cascade displays higher level of phase coherence compared to the one of the helical shell model driven by elongated triads. In the latter case, even though the average phase $\Phi$ appears flattened, a backward flux persists.

### v2

@@ IV.1 Shell spacing dependence
Helical shell models provide a rich framework, in which a set of interacting triads are connected in a way that mirrors the actual interactions among two sequences of helical modes in three dimensional Navier-Stokes turbulence. Here, we consider the phase coherence of these two sequences, as a function of the inter-shell ratio $g$, in order to investigate how differences in synchronizability are related to energy transfer across scales and to compare with the single chain GOY model.
Due to their more complicated structure, helical shell model are less explored in literature. Here, we restrict our analysis to shell models constructed from a single class of interactions, as in Refs. 60, 61, focusing first on those classes that can sustain a forward energy cascade for local interactions, namely SM1, SM2 and SM3. We note that:
  • The separable class SM1 behaves exactly as two uncoupled GOY models. As a result, the intermittency corrections present the same dependence with respect to the inter-shell ratio, and its synchronization properties are equivalent to those of a single chain, as characterized in detail in the first part of this work, showing a the transition from the standard intermittency ($g=2$) to strong intermittency as the inter-shell spacing $g$ is decreased.
@@
  • In stark contrast to all the examples that we have seen earlier, SM3 presents a level of intermittency comparable to SM1 for $g=2$ but the intermittency corrections decrease drastically in the small $g$ limit, and the model tends toward exhibiting monofractal scaling [55].
Those behaviors are presented together in Fig. 12.
Figure 12: Plot of the Scaling exponent $\xi_{p}$of the $p$-th order structure function as a function of $p$, for different Helical shell models corresponding to the classes SM1, SM2, and SM3 (indicated by different colors). The different markers styles represent the two inter-shell spacings $g=2$ and $g=1.1$. The scaling exponents were estimated using the ESS procedure described in Sec.III. The dashed line represents the Kolmogorov $p/3$ scaling.
The time behavior of the SM3 model for the two representative cases with inter-shell ratio $g=1.2$ and $g=2$ is presented in Fig. 13 showing the evolution of the compensated energy spectrum, that can be contrasted to Fig. 10.
In intermittent case ($g=2$), we observe bursty fluctuations around the global scaling $k_{n}^{-2/3}$, less defined but similar to those seen for the GOY model evolution. However no clear signature of instantons can be seen here. It is worth noting that, instantons have been observed in helical shell models when initialized from a zero-background rather than in conditions of a steady state energy cascade. In particular for SM3 class, such structures display an energy spectrum less steep than the Kolmogorov scaling, are reported in Ref. 62.
In contrast, in the non-intermittent case $(g=1.2)$, while some local coherent events, that are carrying energy to smaller scales are still present, they do not seem to propagate across many shells as in the intermittent case since the compensated spectra look more noisy at smaller scales. This suggests that triadic phases lose their ability to self-organize in a way that accommodate strong flux events involving a large portion of the inertial range.
@@
Figure 13: Temporal evolution of compensated energy, $log(k_{n}^{2/3}E_{n}$) across different shells (index $n$ in the $y$-axis) for the two spacing values $g=2$ and $g=1.2$ considered (top to bottom). The plots, corresponding to the SM3 class, illustrate the reduction of coherent structures across scales in the energy cascade as $g$ decreases.
@@
Figure 14: The plots show the decrease in phase coherence observed in the SM3 class as the inter-shell ratio is reduced, for the two representative cases $g=1.2$ (dashed lines) and $g=2$ (solid lines). The top panels displays local coherence showing the normalized PDFs of the global Kuramoto parameter $R$ (left), and the associated average phase $\Phi$ (right). The bottom panels display local coherence, showing the PDFs of the local Kuramoto parameter $R_{n}$ (left) and the associated average phase $\Phi_{n}$ (right) for different shell wave-numbers, as indicated by the colorbar.
@@
The phase coherence in the two regimes is shown in Fig. 14, where normalized PDFs of the global Kuramoto order parameter $R$, and the associated average phase $\Phi$ are presented, for the two inter-shell spacing $g=1.2$ and $g=2$. A clear difference in the level of synchronization is observed, as the system moves from intermittent to non-intermittent behavior, which is manifested by a marked reduction in triadic phase coherence. The monofractal case exhibits a narrow distribution at low $R$ values, in contrast to the broader distribution of $g=2$ case, indicating that incoherent events dominate the dynamics. At the same time, the PDF of the average triadic phase $\Phi$ appears broader as the coherence is lost, even though a clustering around $\pi/2$ persist, which on average, continues to sustain a forward energy flux.
@@
In the bottom panels, we present the PDFs of the parameters quantifying local coherence, as defined in Eq. (10), namely $R_{n}$ and $\Phi_{n}$, for shells within the inertial range. The local level of phase coherence appear to be universal, i.e. independent of the scale. Moreover, the $g=2$ case displays a distribution of local coherence significantly different from that of the GOY model (presented in Fig. 3), also due to the differences in the definition of the local Kuramoto parameters in the two cases, even though both models show comparable levels of intermittency.
@@
Notably, local coherence persists despite the loss of global synchronization as we can see by comparing the PDFs of $R_{n}$ in the two cases $g=2$ and $g=1.2$. When the inter-shell spacing is decreased the position of the maxima shifts slightly toward lower coherence levels, and only highly local coherent events ($R_{n}\simeq 0,8$) are significantly less probable. In contrast, the statistics of both local and global mean triadic phase $\Phi$ show that the phase distribution is more coherent for larger inter-shell spacing. This statistical features are consistent with the observed temporal dynamics, characterized not by structures that travels over many shells but by events that involve neighboring scales.
@@
Regarding forward cascades, we conclude from these results that the Kuramoto parameter, defined using the triadic phases, provides a useful order parameter for studying phase coherence and intermittency in a network of interacting triads. However, it is important to stress that it is not a unique indicator of intermittency. For instance, for the same inter-shell ratio $g=2$ the PDF of the Kuramoto parameter differs between GOY model and the helical model SM3, even though they present very similar levels of intermittency. This observation suggest that the Kuramoto parameter is also sensible to the topological structure of the network of interacting triads. Nonetheless, once the topology is fixed, it proves to be a highly sensible proxy of the modification of anomalous scaling behavior, both when increasing and decreasing the level of intermittency.
@@ V Inverse cascades
Despite the success of shell models in modeling forward cascades, attempts to extend them, in order to reproduce the inverse energy cascade observed in 2D turbulence - simply by changing the dimensionality of the conserved invariants - have generally failed, mainly because of the tendency of the inverse cascade in these models to be overwhelmed by statistical shell-equipartition solutions (see, e.g., Refs. 63, 40). To overcome this issue, different generalization has been proposed, such as shell model on hierarchical trees [64, 65], or models based on spiral chains written for energy [24] or a model only for amplitudes, or one based on a stochastic cascade mechanism [66]. These results suggest that the self-consistently evolving phase equation plays a crucial role in suppressing the inverse cascade.
In this work, we will focus on the two shell model that, to our knowledge, are able to present a genuine inverse energy cascade: the elongated helical shell model proposed by De Pietro, Biferale, and Mailybaev [61] and the model studied by Tom and Ray [50] (see also Ref. 46 for a shell model that shows an inverse cascade driven by resonantly interacting waves in the context of wave turbulence). Our goal here is to analyze the statistical properties of triadic phases and their implication in the inverse cascade solutions produced by these models as a complementary example to the forward cascade case.
Since inverse cascades are boringly non-intermittent and lack coherent bursts, we do not show detailed investigations of the synchronizability of these models. However it is interesting to show the forms of the PDFs of the triadic phases together with the fluxes and the resulting spectra, in order to demonstrate how the triadic phases organize statistically, to drive an inverse cascade even when they are mostly random. We hope that, understanding how triadic phase organization works in such reduced representations as shell models, will provide the key to explaining how the cascade mechanism is sustained in turbulence in general, including the inverse cascade.
@@ V.0.1 Inverse energy cascade in an elongated helical shell model
A particular class of Helical interactions, represented by triads of the type (,+,+)(-,+,+) or (+,,)(+,-,-), i.e. the SM2, is known to drive an inverse cascade for sufficiently elongated triads. The critical elongation, at which the direction of energy transfer changes, (i.e. $p/k=0.278$, assuming $p<k<q$) was first predicted by Waleffe [59], later confirmed with an elongated helical shell model [61], and can apparently, also be justified using arguments based on the conservation of pseudo-invariants [67] in a single triad.
In this section we consider this model, representing a unique example of an inverse cascade driven by non local interactions as our test bed for studying the phase coherence of elongated triads, which drive the inverse cascade process. We recall that for non-local interactions the relevant triadic phases centered at the $n$-th shell can be defined as $u_{n-l}^{s_{1}}u_{n}^{s_{0}}u_{n+m}^{s_{2}}=\chi_{n}e^{i\varphi_{n}}$ and $u_{n-l}^{-s_{1}}u_{n}^{-s_{0}}u_{n+m}^{-s_{2}}=\bar{\chi}_{n}e^{i\bar{\varphi}_{n}}$.
Figure 15: Plot of the energy spectrum as a function of the wave number for a elongated helical shell model (SM2), driven by small-scale forcing (forcing shells indicated in red). The inset plot on the bottom left shows the spectral flux as a function of scale, corresponding to an inverse energy flux in the inertial range. The inset plot on the top right shows the normalized PDFs of triadic phases $\varphi_{n}$, for the shells in the inertial range. The shell number is represented by different colors labeled by the continuous colorbar.
In Figure (15) we present the numerical results for SM2 (elongated) shell model, defined by $s_{1}/s_{0}=-1$ and $s_{2}/s_{0}=+1$, with an inter-shell ratio $g=\frac{1+\sqrt{5}}{2}$ and sufficiently elongated triads, specifically $l=3$ and $m=1$, which satisfy $g^{-l}<0.278$. With this choice of parameters, the introduction of a small scale forcing produces an inverse cascade, as evidenced by the spectral energy flux and the presence of a clear inertial range, with an energy spectrum consistent with the Kolmogorov scaling. The probability distribution functions of non-local triadic phases, presented in the inset (top right), show a higher probability of events with triadic phases equal to $\pi/2$, which corresponds to the choice that maximizes an inverse energy flux. The distribution appears to have a roughly sinusoidal shape and seems to be universal, in the sense that it does not seem to show any scaling with the shell number $n$.
Comparing these PDFs with the forward cascade case, one finds a lower probability of coherent events in the inverse cascade. This weaker phase coherence is even more pronounced at a global level as shown in Figure 2, where the Kuramoto order parameter reveals a relatively incoherent cascade dynamics, characterized by low values of $R$, and a clear flattening of the average phase $\Phi$ distribution.
@@ V.0.2 Inverse energy cascades and equipartition solutions in a local shell model
For the Sabra shell model [17], a detailed phase diagram was proposed in Ref. 68, and a further investigations of the equipartition and the inverse cascade solutions have been carried out by Tom and Ray [50]. Specifically, they considered $g=2$, by setting $a=1$ and imposing energy conservation through the condition c=1bc=-1-b. This causes the second positive definite invariant (in the range 2<b<1-2<b<-1) to have the form $H=\sum_{n}k_{n}^{\alpha}|u_{n}|^{2}\;$ , with a scaling exponent $\alpha=log_{g}\left(\left|\frac{1}{-1-b}\right|\right)$. Notably, the choice of enstrophy conservation, i.e. $\alpha=2$, corresponds to b=5/4b=-5/4 but as we already know this choice fails to drive an inverse cascade because of the issue with the equipartition in shell models. However, they showed that this shell model sustains an inverse energy cascade for values of $b$ in the range $2<b<b_{c}=-1.63$, with a spectral slope that approaches the characteristic scaling of 2D turbulence in the limit $b\to-2$. We have replicated the numerical results presented in Ref. 50, so that we can investigate the role of phase coherence in driving a local inverse cascade.
@@
As shown in Figure 16, the model, when forced at small scales with b=1.8b=-1.8, exhibits a clear inertial range associated with an inverse energy cascade, corresponding to a constant, on average, negative energy flux. The probability distribution of triadic phases, displayed in the top right inset, shows how the triadic phases for different shells organize around the value $\pi/2,$ which maximizes the inverse energy transfer. The distribution follows a sinusoidal shape, i.e odd with respect to $\varphi_{n}\to-\varphi_{n}$, with non universal, scale-dependent behavior, indicating that triadic phase organization is weaker, (as marked by noisier and broader curves) going to larger scales (slower time scales). In contrast (not shown), when the model exhibits energy equipartition (i.e. for $b_{c}<b<-1$) triadic phases appear noisy and randomly distributed.
To quantify synchronization at a global level, providing an overall view of phase coherence between forward and inverse cascades, we report in Figure 2, the global coherence order parameter $R$ and the associated average triadic phase $\Phi$. Notably, we observe that inverse cascades are generally driven by much weaker phase organization. Among the two inverse cascade cases considered, the local inverse cascade displays higher level of phase coherence compared to the one of the helical shell model driven by elongated triads. In the latter case, even though the average phase $\Phi$ appears flattened, a backward flux persists.